OpenStudy (anonymous):

Simplify x^2-36 ------ x^2-12x+36

OpenStudy (anonymous):

= (x-6)(x+6) --------- (x-6)(x-6) = x+6 ---- x-6

OpenStudy (anonymous):

The numerator factors into: x^2 - 36 = (x+6)(x-6) The denominator factors into: x^2 + 12x + 36 = (x+6)(x+6). So we can cancel an (x+6) term and we're left with (x-6)/(x+6) --- Let me know if you have any questions about this!

OpenStudy (hoblos):

(x-6)(x+6) = ---------- (x-6)^2 (x+6) =------- ( x-6)

OpenStudy (hoblos):

LilBarb it is -12 not +12

OpenStudy (anonymous):

My Badd.. o- o

OpenStudy (anonymous):

- *